# The Effectiveness of Curvature-Based Rewiring and the Role of Hyperparameters in GNNs Revisited

### v1

@@ 3 Benchmark datasets have a lack of sufficiently negatively curved edges
Table 1: When an edge is selected by the SDRF algorithm, we verify if this edge satisfies condition 2 of Theorem 4 (Eq. (9)). Additionally, we check if the edge satisfies a softer, but sufficient, version of condition 2 (Eq. (10))
[…]
In our experiments, we make exclusive use of the Stochastic Discrete Ricci Flow (SDRF) [18] algorithm for rewiring, which works in two steps. First, it selects the most negatively curved edge based on the curvature measure. Around this edge, all edges that can lead to three-cycles and four-cycles are considered, and for each candidate edge, the potential improvement of the curvature is computed. The edge to be added is then selected in a stochastic way, regulated by the temperature parameter $\tau$, where the probability is determined by the improvement the edge brings to the curvature of the original edge. Finally, SDRF also allows, at each iteration, the removal of (very) positively curved edges, determined by the threshold curvature value $C^{+}$. Although different algorithms have been proposed (for example BORF where edges are added in batches), we only work with SDRF due to the simplicity of its approach, allowing us to look at the impact of the added edges more clearly.
Looking at the condition from Theorem 4, we now experimentally check if the edges in standard GNN benchmark datasets satisfy the necessary condition 2, which identifies the edge as an ‘oversquashing’ edge. The number of edges to be rewired is chosen from the hyperparameters given in [18]. For each edge selected to be rewired (which is the most negative edge) we compute the upper threshold $\delta_{max}(i,j)=BFc(i,j)+2$. For each of these edges, we then verify if this $\delta_{max}$ determined by the curvature is compatible with the condition
@@
What is the role of different curvature notions?
The question now arises as to what the role is of different curvature measures when rewiring graphs. To analyse this we use SDRF as the standard pre-processing algorithm. We use each discrete curvature notion on graphs discussed in section 2 to rewire the graph and keep track of the selected edges. We then compare, with Balanced Forman curvature as a baseline, the percentage of edges which were selected by SDRF to be rewired. In this process we do not look at the order when the edge was selected, only if the edge was selected. Repeated edges are also considered. The results shown in Table 2 indicate a large discrepancy between the edges selected for rewiring depending on different curvature measures.
Table 2: For each dataset, we rewire the graph 30 times. For each rewiring the number of iterations is randomly generated between the values of the hyperparameter sweep but constant over all curvature measures, while the $C^{+}$ parameter is kept constant, and $\tau$ is taken to be +$\infty$+ ∞. As a reference, we take BFc and for each other curvature measure we compute the % of edges around which is rewired that also were selected with BFc.
Dataset $BFc_{3}/BFc$ $BFc_{mod}/BFc$ $JLc/BFc$ $AFc_{3}/BFc$ $AFc_{4}/BFc$
Texas $76.58\pm 5.36$ $32.92\pm 3.04$ $29.64\pm 5.09$ $16.75\pm 2.27$ $4.31\pm 0.70$
Cornell $39.51\pm 6.24$ $33.24\pm 4.41$ $27.20\pm 4.03$ $17.77\pm 1.91$ $5.37\pm 0.82$
Wisconsin $51.80\pm 6.53$ $32.79\pm 4.23$ $32.28\pm 4.62$ $24.07\pm 2.74$ $6.45\pm 1.40$
Chameleon $12.05\pm 1.25$ $11.88\pm 1.41$ $11.52\pm 1.11$ $2.72\pm 0.74$ $6.75\pm 0.79$
Cora $61.13\pm 5.89$ $52.52\pm 6.56$ $64.99\pm 3.89$ $12.63\pm 1.97$ $6.81\pm 1.36$
Citeseer $54.37\pm 3.49$ $55.90\pm 3.42$ $51.90\pm 3.90$ $10.78\pm 1.55$ $5.73\pm 1.15$
Pubmed $58.22\pm 3.75$ $63.42\pm 3.85$ $56.16\pm 3.24$ $21.17\pm 4.71$ $5.84\pm 1.18$
@@ 4 Hyperparameter dependency of rewiring
To analyse the impact of different curvatures we ran node-classification tasks. For the node classification tasks we used 7 datasets to evaluate the different curvature notions: Texas, Cornell and Wisconsin from WebKB [39], Chameleon [40], Cora [41], Citeseer [42] and Pubmed [43]. For each dataset we only use the largest connected component, extracted with the algorithm provided in https://github.com/jctops/understanding-oversquashing. While this implementation does not strictly adhere to the mathematical definition of connected components for directed graphs (it returns neither the strong nor weak component), it was chosen to maintain consistency in the number of nodes and edges, thus facilitating a fair comparison with other studies. Directed graphs were subsequently transformed to undirected. Our hyperparameter grid is defined as: learning rate: $[0.0001,0.5555]$, number of layers: $\{1,2,3\}$, layer width: $\{16,32,64,128\}$, dropout: $[0.0001,0.5555]$, weight decay: $[0.0001,0.9999]$, $C^{+}$: $[0.2,21.2]$, $\tau$: $[1,500]$ and the max number of iterations is dataset dependent, where we take the lower and upper boundary to be $20\%$ above and below the reported best hyperparameters from [18]. We perform a random-grid search with 1000 iterations for all datasets, except for Pubmed and Chameleon where we performed 500 iterations. Our evaluations follows [27]. Each dataset is split into a development set and a test set with a fixed seed. The development set is then split 100 times into a validation and training set using the same seeds as in [18]. Each hyperparameter configuration is then trained and evaluated on each of the 100 training-validation sets. The reported accuracy is then the mean test accuracy over the 100 validation sets. We train the networks with early stopping, where the patience is 100 epochs of no improvement on the validation set. We use a GCN [2] for all datasets together with the Adam optimizer [44].
@@
Texas Cornell Wisconsin Chameleon Cora Citeseer Pubmed
None $59.95\pm 1.12$ $53.67\pm 0.15$ $54.88\pm 0.49$ $40.96\pm 3.45$ $63.41\pm 14.24$ $58.49\pm 7.40$ $42.21\pm 13.37$
$BFc$ $59.26\pm 0.00$ $53.60\pm 0.25$ $54.05\pm 0.00$ $35.92\pm 2.68$ $28.32\pm 17.70$ $35.54\pm 17.74$ $40.06\pm 9.14$
$BFc_{3}$ $59.26\pm 0.00$ $53.58\pm 0.11$ $54.06\pm 0.00$ $32.31\pm 2.78$ $22.93\pm 11.22$ $29.87\pm 11.09$ $42.61\pm 13.94$
$BFc_{mod}$ $59.26\pm 0.00$ $53.71\pm 0.52$ $55.58\pm 2.13$ $32.42\pm 2.87$ $27.49\pm 9.70$ $45.73\pm 12.01$ $42.96\pm 13.17$
$JLc$ $59.26\pm 0.00$ $53.57\pm 0.00$ $54.06\pm 0.04$ $31.79\pm 2.53$ $26.16\pm 13.47$ $41.82\pm 8.23$ $40.92\pm 13.66$
$AFc_{3}$ $59.63\pm 0.68$ $54.10\pm 1.41$ $55.86\pm 1.63$ $38.39\pm 5.04$ $61.05\pm 14.23$ $60.07\pm 6.37$ $44.19\pm 13.67$
$AFc_{4}$ $59.72\pm 0.83$ $53.64\pm 0.10$ $54.31\pm 0.74$ $31.22\pm 0.68$ $61.03\pm 14.51$ $62.18\pm 5.43$ $43.95\pm 14.23$
Results
@@
Figure 2 show the smoothed distributions (kernel density estimates), together with boxenplots, of the average accuracy obtained per dataset for each rewiring measure. Next to looking at the distributions, we can look at the top $10\%$ of the reported accuracies given by hyperparameters, which is presented in Table 3. In appendix D we also report the maximum value for the mean test accuracy obtained for each dataset and curvature measures. A first observation is that no curvature measure consistently shifts the distribution of the mean test accuracy away from the None distribution (meaning no rewiring) over all datasets. The curvature measure $AFc_{3}$ does provide better performance on datasets such as Cornell and Wisconsin, but it does not do this consistently for the other low homophily datasets Texas and Chameleon. When looking at the average of the top 10% results we see that $AFc$ based rewiring does perform better than other variants while having a smaller standard deviation, but does not advance the performance with respect to no rewiring. It is interesting to note that rewiring can also negatively impacted the performance of a dataset, as seen by the larger spread for the citation networks. Secondly, when comparing the results from Table 1 with the distribution of $BFc$ in Figure 2, we can see that the datasets with more edges that satisfy condition 2b do not perform better than other (e.g. Pubmed). Finally, we can note the similar performance of less-computational intensive curvature measures ($BFc_{3}$ and $AFc_{3}$) in comparison with their more intensive variant ($BFc$ and $AFc_{4}$). This is also interesting to note in the context of Table 2 which showed that the chosen edges to be rewired do not overlap from one curvature to another.
@@
Figure 2: Distribution of mean test accuracy over the sweep of hyperparameters for the different curvature measures and node-classification datasets used. We show boxenplots which first identify the median, then extend boxes outward, each covering half of the remaining data on which outliers (circles) are identifiable. For each dataset we also show the smoothed distribution using kernel density estimates from the seaborn package. The dotted grey line indicates random guessing (1/$\sharp_{classes}$).
@@ 5 Conclusion
While graph rewiring can alleviate structural properties of graphs that cause information bottlenecks, thereby mitigating oversquashing, it also interacts intricately with other factors that influence the performance of a GNN. There is an inherent trade-off between reducing oversquashing and enhancing oversmoothing [29, 21]. Additionally, rewiring can introduce noise, allowing nodes to access information they shouldn’t while improving overall information flow.
Moreover, enabling long-range interactions between nodes is only necessary for tasks that depend on those interactions. This explains why citation networks in this paper, despite meeting theoretical bounds better than other datasets (see Table 1), do not show performance improvements with rewiring. This is in line with previous research indicating that these tasks do not rely on long-range interactions [15]. This dependency is measured by the homophily [35], and Appendix C shows that citation networks have the highest homophily among the benchmark datasets, further strengthening this observation. For the other datasets, the low homophily indicates the necessity of long-range interactions, aligning with the observation that many edges fail to meet theoretical conditions, as shown in Table 1, and therefore explaining the lack of performance gains for these datasets.
Our analysis is further substantiated by examining the role of hyperparameter sweeping when benchmarking curvature-based graph rewiring methods. We found that most performance gain is due to finding an optimal hyperparameter configuration rather than a structural shift in the performance, as illustrated by the comparison between the maximum mean test accuracies (Table 4) and the average of different hyperparameters (Table 3). We argue that future rewiring methods should take into consideration the dependency of their method on hyperparameters, both GNN and rewiring, when evaluating their performance.
@@
The results presented in this manuscript suggest several directions for future investigation while also facing some limitations. While the distributions shown above might be influenced by the hyperparameter grid used in our random search, focusing on the top 10% of results demonstrates that this influence is likely minimal. Our analysis tested the results for the SDRF algorithm from [18], which is the most natural method for curvature-based graph rewiring. It would be interesting to extend this investigation to other rewiring algorithms to validate similar claims. The theoretical analysis of the SDRF bounds indicates that the current bound, $\delta\cdot|\#_{\triangle}|\leq 1$, could potentially be replaced by a more general bound, $\delta\cdot|\#_{\triangle}|\leq R$ with $R>1$, as alluded to in Remark 15 of [18]. However, this adjustment would further weaken the theoretical bounds established in Theorem 4. The curvature distributions provided in the appendix reveal that the edges are not negatively curved enough to create bottlenecks. Future work could involve developing real-world benchmark datasets that do suffer from bottlenecks to test whether the robust performance gains observed in synthetic data can be replicated. Moreover, exploring the possibility of Theorem-aware rewiring, which involves only rewiring edges that meet theoretical conditions, could be beneficial. This approach should be tested on datasets that genuinely suffer from severe information bottlenecks, as indicated earlier.
@@ Appendix B On the implementation of the Balanced Forman curvature
Figure 10: Sample graph provided in Figure. 3 in [18]
During the setup of our experiment, we noticed that the implementation of the Balanced Forman curvature provided by the authors of [18] under https://github.com/jctops/understanding-oversquashing does not match the theoretical definition presented in their paper (Definition 1 in [18]). More precisely, the issue for a given edge $i\sim j$ evolves around the terms corresponding to the four-cycle contributions, i.e. $|\sharp_{\square}^{i}|,\>|\sharp_{\square}^{j}|$ and $\gamma_{max}$. The degree $d_{i}$ and $d_{j}$ of the involved nodes and the number of triangles $|\sharp_{\triangle}(i,j)|$ are calculated correctly. However, even for the sample graph provided in Figure. 3 in [18] (Figure 10 here) the publicly available implementation produces demonstrably wrong results. The four-cycle contribution is computed as illustrated in the code below with $\mathrm{sharp}_{ij}=|\sharp_{\square}^{i}|+|\sharp_{\square}^{j}|$, $\mathrm{lambda}_{ij}=\gamma_{\max}$.
If we consider the edge $0\sim 1$ of the sample graph in Figure 10 and using the definition of the Balanced Forman curvature, Eq. (2) we find $BFc(0,1)=0.10$. In contrast, when using the publicly available code we find $BFc(0,1)=0.08$.
To control that our computation of $BFc$ was correct we implemented with the NetworkX library the set-theoretical definition (evolving around the 1-hop neighbourhoods $S_{1}(i)$ and $S_{1}(j)$ of the involved nodes) provided in Definition 1 in [18]. These are
@@
Additionally, we used Remark 10 from [18] to calculate the four-cycle contribution and in particular $\gamma_{max}$. Our CUDA implementation of BFc was then controlled with the Networkx implementation. Both our implementations are available in our code.
Listing 1: Code snippet from [18] to compute the 4-cycle contribution of the BFc

### v2

@@ 3 Benchmark datasets have a lack of sufficiently negatively curved edges
In our experiments, we make exclusive use of the Stochastic Discrete Ricci Flow (SDRF) (Topping et al., 2021) algorithm for rewiring, which works in two steps. First, it selects the most negatively curved edge based on the curvature measure. Around this edge, all edges that can lead to three-cycles and four-cycles are considered, and for each candidate edge, the potential improvement of the curvature is computed. The edge to be added is then selected in a stochastic way, regulated by the temperature parameter $\tau$, where the probability is determined by the improvement the edge brings to the curvature of the original edge. Finally, SDRF also allows, at each iteration, the removal of (very) positively curved edges, determined by the threshold curvature value $C^{+}$. Although different algorithms have been proposed, we only work with SDRF due to the simplicity of its approach, allowing us to look at the impact of the added edges more clearly.
Looking at the condition from Theorem 4, we now experimentally check if the edges in standard GNN benchmark datasets satisfy the necessary condition 2, which identifies the edge as an ‘oversquashing’ edge. The number of edges to be rewired is chosen from the hyperparameters given in (Topping et al., 2021). For each edge selected to be rewired (which is the most negative edge) we compute the upper threshold $\delta_{max}(i,j)=BFc(i,j)+2$. For each of these edges, we then verify if this $\delta_{max}$ determined by the curvature is compatible with the condition
@@
Table 1: When an edge is selected by the SDRF algorithm, we verify if this edge satisfies condition 2 of Theorem 4 (Eq. equation 9). Additionally, we check if the edge satisfies a softer, but sufficient, version of condition 2 (Eq. equation 10)
Dataset Edges rewired Cond. 2 (%) Cond. 2b (%)
[…]
Chameleon $2441$ $4~{}(0.16\%)$ $141~{}(5.78\%)$
Actor $1000$ ${}(1.1\%)1$ $237~{}(23.70\%)$
[…]
Pubmed $166$ $25~{}(15.06\%)$ $116~{}(69.88\%)$
[…]
@@ 4 Hyperparameter dependency of rewiring
To analyse the impact of different curvatures we ran node-classification tasks. For the node classification tasks we used 9 datasets to evaluate the different curvature notions: Texas, Cornell and Wisconsin from WebKB (University, ), Chameleon and Squirrel (Rozemberczki et al., 2021), Actor (Tang et al., 2009), Cora (McCallum et al., 2000), Citeseer (Sen et al., 2008) and Pubmed (Namata et al., 2012). For graph classification we used MUTAG and PROTEINS (Morris et al., 2020). For each dataset we only use the largest connected component, extracted with the algorithm provided in https://github.com/jctops/understanding-oversquashing. While this implementation does not strictly adhere to the mathematical definition of connected components for directed graphs (it returns neither the strong nor weak component), it was chosen to maintain consistency in the number of nodes and edges, thus facilitating a fair comparison with other studies. Directed graphs were subsequently transformed to undirected.
Our hyperparameter grid is defined as: learning rate: $[0.0001,0.5555]$, number of layers: $\{1,2,3\}$, layer width: $\{16,32,64,128\}$, dropout: $[0.0001,0.5555]$, weight decay: $[0.0001,0.9999]$, $C^{+}$: $[0.2,21.2]$, $\tau$: $[1,500]$ and the max number of iterations is dataset dependent, where we take the lower and upper boundary to be $20\%$ above and below the reported best hyperparameters from (Topping et al., 2021). We perform a random-grid search with 800 iterations for all datasets. Our evaluations follows (Gasteiger et al., 2019). Each dataset is split into a development set and a test set with a fixed seed. The development set is then split 100 times into a validation and training set using the same seeds as in (Topping et al., 2021). Each hyperparameter configuration is then trained and evaluated on each of the 100 training-validation sets. The reported accuracy is then the mean test accuracy over the 100 validation sets. We train the networks with early stopping, where the patience is 100 epochs of no improvement on the validation set. We use a GCN (Kipf & Welling, 2016) for all datasets together with the Adam optimizer (Kingma & Ba, 2014).
Curvature notions
The question now arises as to what the role is of different curvature measures when rewiring graphs. For each dataset, we used each discrete curvature notion on graphs discussed in section 2 to rewire the graph in order to assess possible SDRF as extensively as possible.
Results
Figure 2 show the smoothed distributions (kernel density estimates), together with boxenplots, of the average accuracy obtained per dataset for each rewiring measure. Additionally, we also look at the top $10\%$ of the reported accuracies given by hyperparameters, which is presented in Table 2. To assess the robustness of the distributions we analysed how well the distribution saturates with an increasing number of added iterations. We look at the evolution of the mean and standard deviation as iterations are added, as well as the Wasserstein distance metric between two distributions which include progressively more iterations. We performed this analysis for all datasets and show it in Appendix E. From these results, we see that our sample of iterations is not unrepresentative of the expected performances and that the obtained distributions saturate, indicating that further runs would not majorly impact the performance distribution.
@@ Results
Texas Cornell Wisconsin Chameleon Cora Citeseer
None $59.95\pm 1.15$ $53.66\pm 0.14$ $54.92\pm 0.51$ $40.76\pm 3.52$ $58.83\pm 16.36$ $58.14\pm 7.33$
$BFc$ $59.26\pm 0.00$ $53.61\pm 0.28$ $54.06\pm 0.01$ $34.58\pm 3.19$ $28.39\pm 17.24$ $35.99\pm 13.82$
$BFc_{3}$ $59.26\pm 0.00$ $53.59\pm 0.12$ $54.06\pm 0.01$ $30.93\pm 0.10$ $21.86\pm 08.75$ $30.35\pm 12.03$
$BFc_{mod}$ $59.26\pm 0.00$ $53.68\pm 0.43$ $54.91\pm 1.72$ $31.60\pm 2.04$ $27.73\pm 13.08$ $44.16\pm 12.46$
$JLc$ $59.26\pm 0.00$ $53.57\pm 0.01$ $54.06\pm 0.02$ $30.91\pm 0.02$ $26.83\pm 13.82$ $42.48\pm 7.83$
$AFc_{3}$ $59.58\pm 0.52$ $54.20\pm 1.57$ $56.37\pm 1.60$ $36.93\pm 5.14$ $59.25\pm 14.83$ $60.11\pm 6.30$
$AFc_{4}$ $59.79\pm 0.54$ $53.63\pm 0.10$ $54.60\pm 0.80$ $31.20\pm 0.62$ $58.68\pm 16.10$ $61.67\pm 5.43$
Pubmed Actor Squirrel MUTAG PROTEINS
None $41.99\pm 12.58$ $27.73\pm 0.02$ $36.73\pm 1.96$ $55.34\pm 3.67$ $61.45\pm 1.49$
$BFc$ $39.67\pm 8.30$ $27.73\pm 0.01$ $35.35\pm 1.00$ $54.38\pm 1.71$ $61.36\pm 1.23$
$BFc_{3}$ $40.97\pm 12.01$ $27.73\pm 0.02$ $34.66\pm 0.54$ $54.45\pm 2.25$ $61.16\pm 0.00$
$BFc_{mod}$ $41.23\pm 11.30$ $27.73\pm 0.01$ $34.89\pm 0.85$ $54.56\pm 2.32$ $61.20\pm 0.18$
$JLc$ $39.47\pm 8.97$ $27.73\pm 0.02$ $34.53\pm 0.26$ $54.53\pm 2.86$ $61.20\pm 0.37$
$AFc_{3}$ $42.74\pm 12.91$ $28.00\pm 0.93$ $35.78\pm 1.83$ $54.63\pm 2.91$ $61.22\pm 0.45$
$AFc_{4}$ $41.40\pm 12.65$ $28.14\pm 1.02$ $35.64\pm 1.40$ $54.54\pm 2.40$ $61.27\pm 1.02$
@@
A first observation is that no curvature measure consistently shifts the distribution of the mean test accuracy away from the None distribution (meaning no rewiring) over all datasets. On some datasets, such as Cornell, Wisconsin or Actor we do notice that $AFc$ based rewiring does provide better performance. However, these curvature definitions do not do this consistently for the other low homophily datasets such as Texas, Squirrel or Chameleon. We also see this behaviour when looking at the average of the top 10% results, where $AFc$ based rewiring does perform better than other variants but with a larger standard deviation. The improvement, taking into account the spread of the distributions is therefore not significant with respect to no rewiring. We can also note the similar performance of less-computational intensive curvature measures ($BFc_{3}$ and $AFc_{3}$) in comparison with their more intensive variant ($BFc$ and $AFc_{4}$).
It is also interesting to note that rewiring can also negatively impacted the performance of a dataset, as seen by the larger spread for the citation networks. This could be due to rewiring non-suitable edges which introduces noise in the graph by allowing communication between nodes that shouldn’t. Secondly, when comparing the results from Table 1 with the distribution of $BFc$ in Figure 2, we see that the datasets with more edges that satisfy condition 2b do not perform better than other.
@@
Figure 2: Distribution of mean test accuracy over the sweep of hyperparameters for the different curvature measures and node-classification datasets used. We show boxenplots which first identify the median, then extend boxes outward, each covering half of the remaining data on which outliers (circles) are identifiable. For each dataset we also show the smoothed distribution using kernel density estimates from the seaborn package.
@@ 5 Conclusion
While graph rewiring can alleviate structural properties of graphs that cause information bottlenecks, thereby mitigating oversquashing, it also interacts intricately with other factors that influence the performance of a GNN. There is an inherent trade-off between reducing oversquashing and enhancing oversmoothing (Nguyen et al., 2023; Fesser & Weber, 2023). Additionally, rewiring can introduce noise, allowing nodes to access information they shouldn’t while improving overall information flow.
Moreover, enabling long-range interactions between nodes is only necessary for tasks that depend on those interactions. This explains why citation networks in this paper, despite meeting theoretical bounds better than other datasets (see Table 1), do not show performance improvements with rewiring. This is in line with previous research indicating that these tasks do not rely on long-range interactions (Alon & Yahav, 2021). This dependency is measured by the homophily (Pei et al., 2019), and Appendix C shows that citation networks have the highest homophily among the benchmark datasets, further strengthening this observation. For the other datasets, the low homophily indicates the necessity of long-range interactions, aligning with the observation that many edges fail to meet theoretical conditions, as shown in Table 1, and therefore explaining the lack of performance gains for these datasets.
Our analysis is further substantiated by examining the role of hyperparameter sweeping when benchmarking curvature-based graph rewiring methods. We found that most performance gain is due to finding an optimal hyperparameter configuration rather than a structural shift in the performance, as illustrated by the distribution of performance in Figure 2. We argue that future rewiring methods should take into consideration the dependency of their method on hyperparameters, both GNN and rewiring, when evaluating their performance.
@@
The results presented in this manuscript suggest several directions for future investigation while also facing some limitations. While the distributions shown above might be influenced by the hyperparameter grid used in our random search, focusing on the top 10% of results demonstrates that this influence is likely minimal. Our analysis tested the results for the SDRF algorithm from (Topping et al., 2021), which is the most natural method for curvature-based graph rewiring. It would be interesting to extend this investigation to other rewiring algorithms to validate similar claims. The theoretical analysis of the SDRF bounds indicates that the current bound, $\delta\cdot|\#_{\triangle}|\leq 1$, could potentially be replaced by a more general bound, $\delta\cdot|\#_{\triangle}|\leq R$ with $R>1$, as alluded to in Remark 15 of (Topping et al., 2021). However, this adjustment would further weaken the theoretical bounds established in Theorem 4. The curvature distributions provided in the appendix reveal that the edges are not negatively curved enough to create bottlenecks. Future work could involve developing real-world benchmark datasets that do suffer from bottlenecks to test whether the robust performance gains observed in synthetic data can be replicated. Moreover, exploring the possibility of Theorem-aware rewiring, which involves only rewiring edges that meet theoretical conditions, could be beneficial. This approach should be tested on datasets that genuinely suffer from severe information bottlenecks, as indicated earlier.
@@ Appendix A Curvature distributions
Figure 10: Distribution of curvatures for dataset Actor
Figure 11: Distribution of curvatures for dataset Squirrel
@@ Appendix B On the implementation of the Balanced Forman curvature
Figure 12: Sample graph provided in Figure. 3 in Topping et al. (2021)
During the setup of our experiment, we noticed that the implementation of the Balanced Forman curvature provided by the authors of Topping et al. (2021) under https://github.com/jctops/understanding-oversquashing does not match the theoretical definition presented in their paper (Definition 1 in Topping et al. (2021)). More precisely, the issue for a given edge $i\sim j$ evolves around the terms corresponding to the four-cycle contributions, i.e. $|\sharp_{\square}^{i}|,\>|\sharp_{\square}^{j}|$ and $\gamma_{max}$. The degree $d_{i}$ and $d_{j}$ of the involved nodes and the number of triangles $|\sharp_{\triangle}(i,j)|$ are calculated correctly. However, even for the sample graph provided in Figure. 3 in Topping et al. (2021) (Figure 12 here) the publicly available implementation produces demonstrably wrong results. The four-cycle contribution is computed as illustrated in the code below with $\mathrm{sharp}_{ij}=|\sharp_{\square}^{i}|+|\sharp_{\square}^{j}|$, $\mathrm{lambda}_{ij}=\gamma_{\max}$.
If we consider the edge $0\sim 1$ of the sample graph in Figure 12 and using the definition of the Balanced Forman curvature, Eq. equation 2 we find $BFc(0,1)=0.10$. In contrast, when using the publicly available code we find $BFc(0,1)=0.08$.
To control that our computation of $BFc$ was correct we implemented with the NetworkX library the set-theoretical definition (evolving around the 1-hop neighbourhoods $S_{1}(i)$ and $S_{1}(j)$ of the involved nodes) provided in Definition 1 in Topping et al. (2021). These are
@@
Additionally, we used Remark 10 from Topping et al. (2021) to calculate the four-cycle contribution and in particular $\gamma_{max}$. Our CUDA implementation of BFc was then controlled with the Networkx implementation. Both our implementations are available in our code.
Listing 1: Code snippet from Topping et al. (2021) to compute the 4-cycle contribution of the BFc